OpenStudy (anonymous):

x over 4 = 12 over 16 answer?

OpenStudy (anonymous):

x= 3

OpenStudy (anonymous):

because you cross multiply 16x=48 x=3

OpenStudy (anonymous):

\(\Large \frac{x}{4}=\frac{12}{16} \Leftrightarrow x=\frac{4\times12}{16}=3\)